OpenStudy (anonymous):

determine the linear equation in slope intercept form m=5 (-1,-5)

OpenStudy (anonymous):

i know I use y=Mx+B just not 100% sure where to plug in coordinates

zepdrix (zepdrix):

hmm so i see an m in the question, and an m in the equation you just presented :O hmmmmmmmm

zepdrix (zepdrix):

maybe there's a connection!! c:

OpenStudy (aripotta):

._.

OpenStudy (anonymous):

So far I have Y=5x+B

OpenStudy (aripotta):

i would use the point-slope form first

OpenStudy (aripotta):

then go turn it into slope-intercept

OpenStudy (aripotta):

(y - y1) = m(x - x1)

OpenStudy (aripotta):

or actually, you don't have to do that

OpenStudy (anonymous):

y+5=5(x+1) is that point slope?

OpenStudy (aripotta):

just substitute the x and y into y= 5x + b

OpenStudy (aripotta):

then solve for b

OpenStudy (aripotta):

and then you'll have your answer

OpenStudy (aripotta):

-5 = 5(-1) + b do you know how to solve for b?

OpenStudy (anonymous):

ok then b=0

OpenStudy (aripotta):

yep

OpenStudy (aripotta):

so y = 5x

OpenStudy (aripotta):

do you understand?

OpenStudy (anonymous):

what does the B mean?

OpenStudy (aripotta):

b is the y-intercept

OpenStudy (anonymous):

is that my start point

OpenStudy (aripotta):

yea

OpenStudy (aripotta):

so you get why it's y = 5x?

OpenStudy (anonymous):

yup